OpenStudy (anonymous):

The base of a triangle is 3 cm longer than its altitude. The area of the triangle is 35cm^2. Find the altitude.

OpenStudy (freckles):

Area of triangle=b*h/2 you have b=3+h the area is 35 We have 35=b*h/2 or you can say 70=b*h

OpenStudy (freckles):

Replace b with 3+h in 70=b*h